OpenStudy (chris215):

Find the number a such that the line x = a divides the region bounded by the curves y = x, y = 0, and x = 4 into two regions with equal area.

OpenStudy (plainntall):

I find it helps to graph what we are talking about. 

OpenStudy (sshayer):

area fromx= 0 to a =area fromx= a to 4

OpenStudy (sshayer):

\[\int\limits_{o}^{a}y dx=\int\limits_{a}^{4}y dx\] \[\int\limits_{0}^{a}x dx=\int\limits_{a}^{4}x dx\] complete it

OpenStudy (chris215):

I got a=2sqrt(2) =2.828

OpenStudy (mathmale):

How would you go about checking y our result? Such a check would be pretty straightforward.

OpenStudy (sshayer):

\[a=2\sqrt{2}\] is exact. but decimal is approximate

OpenStudy (mathmale):

Yes. Now, if you use this a as the upper limit of your first definite integral, and as the lower limit of your second, you could calculate a numerical answer (an area, by the way) for each integral. Are the 2 resulting def. ints. equal to one another? I agree that 2Sqrt(2) is "exact" and as such is a better answer than "2.828."

OpenStudy (chris215):

the directions told me to give my answer to 3 decimal places so

OpenStudy (chris215):

thank you!

OpenStudy (sshayer):

yw